Let f(x) = 2x – 1.9(x) = 3x, and h(x) = x2 + 1. Compute the following: (h • (gof)(x)
Answer #1

g(f(x)) = g(2x-1) = 3(2x-1)

= 6x -3

h(g(f(x))) = h(6x-3)

=(6x-3)2 + 1 = 36x2-36x+9+1 = 36x2​​​​​​​-36x+10